OpenStudy (anonymous):

find the coefficient of x^4 in the expansion of (3x-1)^11

OpenStudy (anonymous):

same as before \[\dbinom{11}{4}(3x)^4(-1)^7\]

OpenStudy (anonymous):

and what does the 11 over 4 part mean again?

OpenStudy (anonymous):

\[3^4=81\] \[(-1)^7=-1\]

OpenStudy (anonymous):

do you know how to find \[\dbinom{11}{4}\]? if not i will be happy to show you.

OpenStudy (anonymous):

it is called "eleven choose 4" the number of ways to choose 4 items out of 11.

OpenStudy (anonymous):

show please!! also why did u just do 3^4 isnt ther an x in there like 3x^4?

OpenStudy (anonymous):

\[(3x)^4=3^4x^4\] and since you were asked for the coefficient that is the number you need to compute.

OpenStudy (anonymous):

it is the whole thing raised to the power of 4, not just the x. that is it is \[(3x)^4\] not \[3x^4\]

OpenStudy (anonymous):

so there for it is -81x^4

OpenStudy (anonymous):

if i times it by the -1 like i was supposed to?

OpenStudy (anonymous):

now to compute \[\dbinom{11}{4}\] yes there is a -81

OpenStudy (anonymous):

but you have to multiply by \[\dbinom{11}{4}\] which is easy enough to compute. i can show you step by step if you like

OpenStudy (anonymous):

thatd be great!

OpenStudy (anonymous):

ok first of all a formula, although you don't really use it. the formula is \[\dbinom{n}{k}=\frac{n!}{k! (n-k)!}\] here n = 11, k = 4 and n-k=7 make a fraction. in the numerator put 4 numbers starting at 11 and counting down. in the denominator you put 4! to get \[\dbinom{11}{4}=\frac{11\times 10 \times 9\times 8}{4 \times 3\times 2}\]

OpenStudy (anonymous):

is my final answer -26730x^4?

OpenStudy (anonymous):

now since this is a whole number , cancel first and multiply last! \[\frac{11 \times 10\times 9 \times 8}{4\times 3\times 2}={11\times 10\times 3}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\dbinom{11}{4}=330\] \[330\times 81 \times -1=-26730\] yes you got it!

OpenStudy (anonymous):

thankyou so much!!!!!!! your awesome!

OpenStudy (anonymous):

can i ask u about another??!

OpenStudy (anonymous):

no problem. try \[\dbinom{10}{3}\] and convince yourself that it is the same as \[\dbinom{10}{7}\]because 3+7=10!

OpenStudy (anonymous):

sure ask away.

OpenStudy (anonymous):

find the coefficient of x^7 in the expansion of (2x-5)^9

OpenStudy (anonymous):

ok same idea. here n = 9, k = 7, n-k=3 so the term with \[x^7\] will look like \[\dbinom{9}{7}(2x)^7(-5)^3\]

OpenStudy (anonymous):

oops typo sorry

OpenStudy (anonymous):

n-k=2!

OpenStudy (anonymous):

my mistake. it will be \[\dbinom{9}{7}(2x)^7(-5)^2\]

OpenStudy (anonymous):

exponents have to add up to 9. would you like to try it?

OpenStudy (anonymous):

so what would (2x)^7 be?

OpenStudy (anonymous):

\[(2x)^7=2^7x^7\] so you will need to compute \[2^7\]

OpenStudy (anonymous):

thats is easy, as is \[(-5)^2\]

OpenStudy (anonymous):

your real job is to compute \[\dbinom{9}{7}\]

OpenStudy (anonymous):

yes how do u do the 9 over 7 thging again?

OpenStudy (anonymous):

\[\dbinom{n}{k}=\frac{n!}{k! (n-k)!}\]

OpenStudy (anonymous):

here n = 9, k = 7 and n-k=2

OpenStudy (anonymous):

so 9! over 7! times 2!

OpenStudy (anonymous):

right. but don't forget to cancel away first because the entire denominator will cancel

OpenStudy (anonymous):

the answer is 36?

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

so 36 times 128 times 25

OpenStudy (anonymous):

so 25200x^7?

OpenStudy (anonymous):

now i show you the easy way. first of all 7+2=9 so it is easier to compute \[\dbinom{9}{2}\] so we work as before. make a fraction. in the numerator put to numbers starting at 9 in the denominator put 2. we get the answer right away. \[\dbinom{9}{2}=\frac{9\times 8}{2}=9\times 4=36\]

OpenStudy (anonymous):

good!!

OpenStudy (anonymous):

yes, 36 times 128 times 25 is it.

OpenStudy (anonymous):

oh wait i got a different number than you. i got 115200

OpenStudy (anonymous):

maybe i put it in wrong.

OpenStudy (anonymous):

no i think i am right.

OpenStudy (anonymous):

nah i put it in wrong!

OpenStudy (anonymous):

whew i was scared but it is late.

OpenStudy (anonymous):

is that enough of this? or are there more?

OpenStudy (anonymous):

you have to go?

OpenStudy (anonymous):

i can help you with another if you like.

OpenStudy (anonymous):

one question quick.. the fianl answer is 115200x^7 rights?

OpenStudy (anonymous):

final answer yes.

OpenStudy (anonymous):

nah its okay i will let u go! u were a great help! quick question .. ur suyper smart!! how old are u?!

OpenStudy (anonymous):

old as black pepper. have fun, and don't forget to convince yourself that \[\dbinom{10}{6}=\dbinom{10}{4}\]because 10=4+6

OpenStudy (anonymous):

awesome thanks!! haha